# On rate-optimal partitioning classification from observable and from privatised data

### v2

@@ 4 Proofs
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $\leq\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}\Delta(x)^{2}\right)\mu(dx)$
$+\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}\Delta(x)^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx).$ (27)
Similarly to (4.1), the margin condition implies
$\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}(P_{(1)}(x)-P_{(2)}(x))^{2}\right)\mu(dx)$
$=\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G^{*}(ds)$
$=O\left((nh^{d_{a}})^{-(\gamma+1)/2}\right),$ (28)
where $a=(nh^{d_{a}})/(8M^{2})$. As (22), the combined margin and density condition yields that
$\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int_{S_{a}}\exp\left(-\frac{nh^{d_{a}}}{8M^{2}}(P_{(1)}(x)-P_{(2)}(x))^{2}f_{h}(x)\right)\frac{1}{\sqrt{f_{h}(x)}}\mu_{a}(dx)$
$=\frac{4M^{2}}{\sqrt{nh^{d_{a}}}}\int_{0}^{\infty}e^{-as^{2}}G_{\sqrt{f_{h}}}(ds)$
$=O\left((nh^{d_{a}})^{-(\gamma(\sqrt{f})+1)/2}\right).$ (29)
Thus,
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh^{d_{a}})^{-(\gamma(\sqrt{f})+1)/2}\right)+O\left((nh^{d_{a}})^{-(\gamma+1)/2}\right).$ (30)
@@
$J_{n,k,l}^{(2)}$ $=\sum_{j}\int_{A_{h,j}}R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq R_{l}^{*}(x)/(2M)\}}\mu(dx)$
$=\sum_{j}\int_{A_{h,j}}R_{l}^{*}(x){\mathbb{I}}_{\{|\nu_{k}(A_{h,j})-\mu(A_{h,j})P_{k}(x)|\geq\mu(A_{h,j})R_{l}^{*}(x)/(2M)\}}|m(x)|\mu(dx).$
@@
$J_{n,k,l}^{(2)}$ $\leq\sum_{j}\int_{A_{h,j}}{\mathbb{I}}_{\{C\sqrt{d}h\geq R_{l}^{*}(x)/(2M)\}}R_{l}^{*}(x)\mu(dx)\leq c^{*}(2C\sqrt{d}Mh)^{1+\gamma},$ (31)
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=M^{2}O\left((Mh)^{1+\gamma}\right)=M^{3+\gamma}O\left(h^{1+\gamma}\right).$ (32)
@@
$I_{n}+{\mathbb{E}}\{J_{n}\}$ $=O\left(h^{1+\gamma}\right)+O\left(\frac{1}{(nh^{d_{a}})^{(\min\{1,\gamma,\gamma(\sqrt{f^{*}})\}+1)/2}}\right).$ (33)
@@
${\mathbb{E}}\{K_{n}\}$ =$=\int{\mathbb{P}}\left\{|{{\widehat{m}}^{\prime}}_{n}(x)|\geq|m(x)|/3\right\}|m(x)|\mu(dx)$
@@
$\leq 6\max_{z}\frac{z}{1+z^{3}}\sum_{j}\frac{c\varrho_{\varepsilon}\sigma_{Z}}{\sigma_{\varepsilon}^{3}n}$
$=\frac{\tilde{c}\cdot\sigma_{Z}}{nh^{d}}$
@@
$\leq 2\sum_{j}\int_{A_{h,j}}\frac{3\sigma_{Z}}{\sqrt{n}\mu(A_{h,j})|m(x)|}\exp\big{(}-\frac{n\mu(A_{h,j})^{2}m(x)^{2}}{18\sigma_{Z}^{2}}\bigg{)}|m(x)|\mu(dx).$
@@
$+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$+2\int\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$2\int\frac{3}{\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nh^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu(dx)$ $=O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma+1)/2}\right).$
and by the combined margin and density condition
@@
$=O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma(f)+1)/2}\right).$
Summary:
${\mathbb{E}}(L(\widetilde{g}_{n}))-L^{*}$
$=O\left(h^{1+\gamma}\right)+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma+1)/2}\right)+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma(f)+1)/2}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
@@
$J_{n,k,l}^{(1)}$ $=\int R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq R_{l}^{*}(x)/(3M)\}\mu(dx),$
@@
$J_{n,k,l}^{(2)}$ $=\int R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq R_{l}^{*}(x)/(3M)\}}\mu(dx),$
@@
$Q_{n,k}(x)=\frac{\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}}{\mu(A_{n,j})},\quad\mbox{ if }x\in A_{n,j}.$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh^{d_{a}})^{-(\gamma(\sqrt{f^{*}})+1)/2}\right)+O\left((nh^{d_{a}})^{-(\gamma+1)/2}\right),$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=O\left(h^{1+\gamma}\right).$
@@
$R_{l}^{*}(x){\mathbb{I}}_{\{l\neq g^{*}(x)\}}{\mathbb{P}}\left\{\left|\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}\right|\geq\mu(A_{h,j})R_{l}^{*}(x)/(3M)\right\}$
$\leq\frac{3MR_{l}^{*}(x)}{\sqrt{n}\mu(A_{h,j})R_{l}^{*}(x)}{\mathbb{I}}_{\{l\neq g^{*}(x)\}}\exp\left(-\frac{n\mu(A_{h,j})^{2}R_{l}^{*}(x)^{2}/(3M)^{2}}{4\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right)$
$\leq\frac{3M}{\sqrt{n}\mu(A_{h,j})}{\mathbb{I}}_{\{l\neq g^{*}(x)\}}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}/(3M)^{2}}{4\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
For $h$ small enough, this together with the arguments in (20), (4.1), (22), (4.3) and (4.3) imply that
@@
$\leq 2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}<\mu(A_{h,j})\}}\int_{A_{h,j}}\frac{3M}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}/(3M)^{2}}{4\sigma_{Z}^{2}}\right)\mu(dx)$
$+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{3M}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu_{a}(A_{h,j})^{2}\Delta(x)^{2}/(3M)^{2}}{4\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(3)}=O\bigg{(}\Big{(}\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\Big{)}^{(\gamma+1)/2}\bigg{)}+O\bigg{(}\Big{(}\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\Big{)}^{(\gamma(f)+1)/2}\bigg{)}+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$

### v3

@@ 5 Proofs
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right)+O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right)+O\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}.$ (31)
@@
$J_{n,k,l}^{(2)}$ $=\sum_{j}\int_{A_{h,j}}\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/4\}}\mu(dx)$
$=\sum_{j}\int_{A_{h,j}}\Delta_{l}(x){\mathbb{I}}_{\{|\nu_{k}(A_{h,j})-\mu(A_{h,j})P_{k}(x)|\geq\mu(A_{h,j})\Delta_{l}(x)/4\}}|m(x)|\mu(dx).$
@@
$J_{n,k,l}^{(2)}$ $\leq\sum_{j}\int_{A_{h,j}}{\mathbb{I}}_{\{C\sqrt{d}h\geq\Delta_{l}(x)/4\}}\Delta_{l}(x)\mu(dx)\leq c^{*}(4C\sqrt{d}h)^{1+\gamma},$ (32)
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=M^{2}O\left(h_{n}^{1+\gamma}\right).$ (33)
@@
$I_{n}+{\mathbb{E}}\{J_{n}\}$ $=O\left(h_{n}^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right).$
@@
${\mathbb{E}}\{K_{n}\}$ $=\int{\mathbb{P}}\left\{|{{\widehat{m}}^{\prime}}_{n}(x)|\geq|m(x)|/3\right\}|m(x)|\mu(dx)$
@@
$\leq 6\max_{z}\frac{z}{1+z^{3}}\sum_{j}\frac{c\varrho_{\varepsilon}\sigma_{Z}}{\sigma_{\varepsilon}^{3}n}=\frac{\tilde{c}\sigma_{Z}}{nh^{d}}$
@@
$\leq 2\sum_{j}\int_{A_{h,j}}\frac{3\sigma_{Z}}{\sqrt{n}\mu(A_{h,j})|m(x)|}\exp\bigg{(}-\frac{n\mu(A_{h,j})^{2}m(x)^{2}}{18\sigma_{Z}^{2}}\bigg{)}|m(x)|\mu(dx).$
@@
$\quad+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$\quad+2\int\frac{3}{f_{h}(x)\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nf_{h}(x)^{2}h^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$2\int\frac{3}{\sqrt{nh^{2d_{a}}/\sigma_{Z}^{2}}}\exp\left(-\frac{nh^{2d_{a}}m(x)^{2}}{18\sigma_{Z}^{2}}\right)\mu(dx)$ $=O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(\gamma+1)/2}\right).$
and by the modified combined margin and density condition
@@
$=O\left(\left(\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\right)^{(\gamma_{2}+1)/2}\right).$
In conclusion, we have
${\mathbb{E}}(L(\widetilde{g}_{n}))-L^{*}=O\left(h^{1+\gamma}\right)+O\left(\bigg{(}\frac{1}{nh_{n}^{d_{a}}}\bigg{)}^{(\min\{1,\gamma,\gamma_{1})\}+1)/2}\right)$
$\quad+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma+1)/2}\right)+O\left(\left(\frac{\sigma_{Z}^{2}}{nh^{2d_{a}}}\right)^{(\gamma_{2}+1)/2}\right)+O\left(\frac{\sigma_{Z}}{nh^{d}}\right).$
@@
$J_{n,k,l}^{(1)}$ $=\int\Delta_{l}(x){\mathbb{P}}\{|P_{n,k}(x)-{\mathbb{E}}\{P_{n,k}(x)\}|\geq\Delta_{l}(x)/6\}\mu(dx),$
@@
$J_{n,k,l}^{(2)}$ $=\int\Delta_{l}(x){\mathbb{I}}_{\{|{\mathbb{E}}\{P_{n,k}(x)\}-P_{k}(x)|\geq\Delta_{l}(x)/6\}}\mu(dx),$
@@
$Q_{n,k}(x)=\frac{\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}}{\mu(A_{h,j})},\quad\mbox{ if }x\in A_{h,j}.$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(1)}$ $=O\left((nh_{n}^{d_{a}})^{-(\gamma_{1}+1)/2}\right)+O\left((nh_{n}^{d_{a}})^{-(\gamma+1)/2}\right),$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(2)}$ $=O\left(h_{n}^{1+\gamma}\right).$
@@
$\Delta_{l}(x){\mathbb{P}}\left\{\left|\frac{\sigma_{Z}}{n}\sum_{i=1}^{n}\epsilon_{i,j,k}\right|\geq\mu(A_{h,j})\Delta_{l}(x)/6\right\}$
$\leq\frac{6\Delta_{l}(x)}{\sqrt{n}\mu(A_{h,j})\Delta_{l}(x)}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta_{l}(x)^{2}/6^{2}}{4\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right)$
$\leq\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$
For $h$ small enough, this together with the arguments in (21), (5.1), (23), (5.3) and (5.3) imply
@@
$\leq 2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}<\mu(A_{h,j})\}}\int_{A_{h,j}}\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)\mu(dx)$
$\quad+2\sum_{j}{\mathbb{I}}_{\{h^{d_{a}}\geq\mu(A_{h,j})>0\}}\int_{A_{h,j}}\frac{6}{\sqrt{n}\mu(A_{h,j})}\exp\left(-\frac{n\mu_{a}(A_{h,j})^{2}\Delta(x)^{2}}{144\sigma_{Z}^{2}}\right)\mu_{a}(dx)$
@@
$\sum_{k=1}^{M}\sum_{l=1}^{M}J_{n,k,l}^{(3)}=O\bigg{(}\bigg{(}\frac{\sigma_{Z}^{2}}{nh_{n}^{2d_{a}}}\bigg{)}^{(1+\min(\gamma,\gamma_{2}))/2}\bigg{)}+O\left(\frac{\sigma_{Z}}{nh_{n}^{d}}\right).$